A newspaper article reported that people spend a mean of 6.5 hours per day watching TV, with a standard deviation of 1.7 hours. A psychologist would like to conduct interviews with the 15% of the population who spend the most time watching TV. She assumes that the daily time people spend watching TV is normally distributed. At least how many hours of daily TV watching is necessary for a person to be eligible for the interview. Carry your intermediate comutations to a least four decimal places. Round your answer to at least one decimal place.
Thank you.

Answer by stanbon(75887)

A newspaper article reported that people spend a mean of 6.5 hours per day watching TV, with a standard deviation of 1.7 hours.
A psychologist would like to conduct interviews with the 15% of the population who spend the most time watching TV.
She assumes that the daily time people spend watching TV is normally distributed. At least how many hours of daily TV watching is necessary for a person to be eligible for the interview. Carry your intermediate computations to a least four decimal places. Round your answer to at least one decimal place.
---------------------------------
mean = 6.5 hrs ; s = 1.7 hrs
Draw the normal curve picture..
Put a mark on the horizontal access that has 15% of the population to the right.
----------------------
Using your chart or calculator find the z value of that point:
z = InvNorm(0.85) = 1.0364..
Now find the raw score corresponding to that z-value, as follows:
z = (x-u)/s
1.0364 = (x-6.5)/1.7
x-6.5 = 1.7*1.0364
x = 8.2619... hrs.
People who watch more that 8.2619 hrs of TV per day may be interviewed.
==================
Cheers,
Stan H.
==========================================
Cheers,
Stan H.